1) A quartz piezoelectric crystal having a thickness of 2 mm and a voltage sensitivity of 0.055 V · m/N is subjected to a pressure of 200 psi. Calculate the voltage output.

2) A Hall-effect transducer is used for the measurement of a magnetic field of 5000 G. A 2-mm slab of bismuth is used with a current of 3 A. Calculate the voltage output of the device.

3)In a Michelson interferometer, 100 fringes cross the field of view when the movable mirror is displaced by 0.02948mm.Calculate the wavelength of monochromatic source.

4) A certain microphone has an open circuit sensitivity of ?80 dB referenced to 1 V for sound-pressure excitation of 1dyn/cm2.Calculate the voltage output when exposed to a sound field of (a) 100 dB and (b) 50 dB
Answer #1

1)ANSWER

Output voltage of piezoelectric crystal can be calculated by using formula :

V = P×g× t

where P is the pressure applied in N/(sq.m );g=sensitivity of material;t:thickness of material and V is output voltage

V=0.055

Given data:

p=200 psi =200*6894.76 where 1psi(pound force per square inch)=6894.76

=1378952.0 N/sq.m  

now ,substituting we've ,

V=1378952*0.055*2*10^(-3)

=151.68V